Circular No. 6852

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980326 P. J. Groot and P. M. Vreeswijk, University of Amsterdam; E. Pian

and E. Palazzi, CNR Bologna; C. Tinney, Anglo-Australian Observatory; M. Keane, National Optical Astronomy Observatory; and J. F. Gonzalez, European Southern Observatory (ESO), report on behalf of a larger collaboration: “Comparison of optical R-band observations of the error box of GRB 980326 (IAUC 6851) taken at the Anglo-Australian Telescope (AAT) on Mar. 27.42 UT, the Cerro Tololo 4-m telescope on Mar. 28.04, and the ESO 3.5-m New Technology Telescope on Mar. 28.01 and 28.17 shows a clearly variable object located at α = 8^{h}36^{m}34^{s}.28, δ = −18^{o}51′23′′.9 (equinox 2000.0). Differential photometry with respect to three local stars shows decreases in brightness between Mar. 27.42 and 28.01 of 1.2(1) mag; between Mar. 27.42 and 28.04 of 1.6(1) mag; and between Mar. 27.42 and 28.17 of 1.7(1) mag. The three comparison stars have position end figures and magnitudes (determined with respect to Landolt SA 98 stars 626, 624, 634, 590) as follows: 35^{s}.41, 51′41′′.6, R = 18.4(1); 31^{s}.00, 51′17′′.3, 21.7(1); 34^{s}.03, 50′53′′.2, 20.3(1). At the time of the AAT observation (Mar. 27.42), the magnitude of the variable was R = 21.7, equal to that of the second comparison star.”

SUPERNOVA 1995ad IN NGC 2139 J. Broughton, Reedy Creek, Queensland, reports the following predis-

covery position and magnitude of SN 1995ad (cf. IAUC 6239, 6242) from a CCD frame taken with a 0.25-m f/6.3 reflector on 1995 Sept. 22.79 UT: α = 6^{h}01^{m}06^{s}.11, δ = −23^{o}40′29′′.1 (equinox 2000.0), R = 15.7.

SUPERNOVA 1998S IN NGC 3877 Visual magnitude estimates: Mar. 23.39 UT, 12.0 (A. Mizser, Bu-

dapest, Hungary); 23.860, 12.3 (A. Diepvens, Balen, Belgium); 24.09, 11.8 (M. S. Tejera, Las Palmas, Canary Is., Spain); 24.8, 12.0 (L. Kiss, Szeged, Hungary); 24.98, 11.7 (J. Carvajal, Madrid, Spain); 25.10, 11.7 (W. Dillon, Damon, TX); 25.32, 12.0 (R. Keen, Mt. Thorodin, CO); 25.881, 12.1 (A. Pereira, Cabo da Roca, Portugal); 26.774, 12.0 (M. Plsek, Lelekovice, Czech Republic); 26.79, 12.0 (A. Baransky, Kiev, Ukraine); 26.95, 11.9 (Carvajal).

1998 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green